OpenStudy (cenaida):

What are the coordinates of the vertex of the parabola given by the equation: y = x2 - 2x - 2 ? A. (0, -2) B. (1, -3) C. (4, 6) D. (-2, 6)

OpenStudy (cenaida):

@SACAPUNTAS

OpenStudy (anonymous):

Oh, this is like the one we just did. Remember the formula for the vertex? :)

OpenStudy (cenaida):

im sorry I don't really remember :/

OpenStudy (anonymous):

This literally starts exactly the same as the last question we did. :D Go look at it again. I'll wait.

OpenStudy (cenaida):

the vertex formula is −b/2a

OpenStudy (anonymous):

Yes. So start by plugging into that. That gives us the x coordinate of the vertex. Then you can find y via the equation itself, \(y = x^2 - 2x - 2\)

OpenStudy (anonymous):

Oh, I should have asked, you know where the a, b, c, etc. are coming from, right?

OpenStudy (cenaida):

no..

OpenStudy (anonymous):

Oh! I'm sorry. That makes this a lot less clear. It's the standard form of the equation: \(ax^2 + bx + c = 0\) As long as it's in that form there are a lot of useful formulas you can use, and they'll all talk about a, b and c. So in this case, the equation is \(y = x^2 - 2x - 2\) which means that a = 1, b = -2, and c = -2

OpenStudy (cenaida):

hmm is the answer D?

OpenStudy (anonymous):

Let's check! \(\frac{-b}{2a} = \frac{-(-2)}{2(1)} = \frac{2}{2} = 1\) So the x value is 1. I'm not even going to bother to find y because only one of the choices has an x value of 1.

OpenStudy (cenaida):

y is -3 which makes the answer B

OpenStudy (anonymous):

Yes. :)

OpenStudy (cenaida):

thanks!!